A student driving to class notices that her daily commute takes anywhere between 10 and 15 minutes every day. Let’s assume that the daily commute has a Uniform probability distribution, that is any time between 10 and 15 minutes is equally likely to occur. What is P(12< X <14)?
Answer #1

Solution :

Given that,

a = 10

b = 15

P(c < x < d) = (d - c) / (b - a)

P(12 < x < 14) = (14 - 12) / (15 - 10)

= 2 / 5 = 0.4

P(12 < x < 14) = 0.4